x^(1/2)-3^(1/4)+2=0
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

x^(1/2)-3^(1/4)+2=0
------------------
This is a quadratic; the variable is x^(1/4)
---------------------------------
Let w = x^(1/4)
----------------------
w^2 - 3w + 2 = 0
(w-2)(w-1) = 0
w = 2 or w = 1
--------------------------
x^(1/4) = 2 or x^(1/4) = 1
x = 16 or x = 1
================================
Cheers,
Stan H.

Answer by Alan3354(69443)

x^(1/2)-3^(1/4)+2=0
x^(1/2) = -2 + 3^(1/4)
Square both sides
x = 4 + sqrt(3) - 2*3^(1/4)
x = ~ 4 + 1.73205 - 2*1.316017
x = ~ 3.0999